OpenStudy (anonymous):

The circulation of a newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation?

OpenStudy (anonymous):

\[percent~decrease=\frac{ difference }{ original } \times 100 =\frac{ 3200-2464 }{ 3200 } \times 100\]